# Discovering Preference Optimization Algorithms with and for Large Language Models

### v1

@@ Abstract
Offline preference optimization is a key method for enhancing and controlling the quality of Large Language Model (LLM) outputs. Typically, preference optimization is approached as an offline supervised learning task using manually-crafted convex loss functions. While these methods are based on theoretical insights, they are inherently constrained by human creativity, so the large search space of possible loss functions remains under explored. We address this by performing LLM-driven objective discovery to automatically discover new state-of-the-art preference optimization algorithms without (expert) human intervention. Specifically, we iteratively prompt an LLM to propose and implement new preference optimization loss functions based on previously-evaluated performance metrics. This process leads to the discovery of previously-unknown and performant preference optimization algorithms. The best performing of these we call Discovered Preference Optimization (DiscoPOP) [^1] Code: https://github.com/luchris429/DiscoPOP., a novel algorithm that adaptively blends logistic and exponential losses. Experiments demonstrate the state-of-the-art performance of DiscoPOP and its successful transfer to held-out tasks.
@@ 1 Introduction
In this work, we aim to address these limitations by performing LLM-driven discovery in order to automatically generate new state-of-the-art preference optimization algorithms without continual expert human intervention in the development process. While previous works (Ma et al., 2023; Yu et al., 2023) have used LLMs to design environment-specific RL reward functions, we discover general-purpose objective functions which can be used across various preference optimization tasks. More specifically, we iteratively prompt an LLM to propose new preference optimization loss functions and evaluate them, with the previously proposed loss functions and their task performance metric (in our case, MT-Bench scores (Zheng et al., 2024)) as in-context examples. After performing this automatic discovery process, we catalog high-performing loss functions and introduce a particularly strong one we call Discovered Preference Optimization (DiscoPOP), a new algorithm. To ensure robustness beyond MT-Bench, we validate DiscoPOP using AlapacaEval 2.0 (Dubois et al., 2024), showing an improvement in win rates against GPT-4 from DPO $(11.23\%\rightarrow 13.21\%)$. Additionally, in separate, held-out, tasks such as summarization and controlled generation, models trained with the DiscoPOP loss outperform or perform competitively with existing preference optimization algorithms.
@@
Figure 1: Left. Conceptual illustration of LLM-driven discovery of objective functions. We prompt an LLM to output new code-level implementations of offline preference optimization losses $\mathbb{E}_{(y_{w},y_{l},x)\sim\mathcal{D}}\left[f\left(\beta\rho\right)\right]$ as a function of the policy ($\pi_{\theta}$) and reference model’s ($\pi_{\text{ref}}$) likelihoods of the chosen ($y_{w}$) and rejected ($y_{l}$) completions. Afterward, we run an inner loop training procedure and evaluate the resulting model on MT-Bench. The corresponding performance is fed back to the language model, and we query it for the next candidate. Right. Performance of discovered objective functions on Alpaca Eval.
@@ 2 Background
Preference Optimization. Consider a pre-trained language model policy $\pi_{\theta}$ and a dataset $\mathcal{D}=\{(x^{i},y_{w}^{i},y_{l}^{i})\}_{i=1}^{N}$ consisting of prompts $x$ and preference-ranked completions $y_{w}$ and $y_{l}$. In this dataset, a human rater prefers $y_{w}$ over $y_{l}$, denoted as $y_{w}\succ y_{l}$. The task is to align $\pi_{\theta}$ with the human values implicit in these preferences. Canonically, this has been achieved through reinforcement learning from human feedback (Christiano et al., 2017, RLHF), an approach that proceeds in two phases: First, a reward modeling stage that learns a parameterized reward model $r_{\phi}$. By assuming a Bradley-Terry model (Bradley and Terry, 1952) of preferences, the probability of the data can be expressed as $P(y_{w}\succ y_{l})=\exp r_{\phi}({y_{w},x})/(\exp r_{\phi}({y_{w},x})+\exp r_{\phi}({y_{l},x}))$, and subsequently simply optimized over $\phi$ through the maximum likelihood principle. The second stage of policy optimization employs a reinforcement learning algorithm to train the language model against the learned reward. Usually, a KL penalty is introduced between the model and the pre-RL reference policy $\pi_{ref}$ (Jaques et al., 2019; Stiennon et al., 2020) to prevent over-optimization and straying too far from the original policy, resulting in the final objective:
@@
Despite success in frontier models (Anthropic, 2023; Gemini-Team, 2023), deep RL has many implementation (Engstrom et al., 2019) and training challenges (Sutton, 1984; Razin et al., 2023) that hinder its adoption. In order to simplify the whole process, direct preference optimization (Rafailov et al., 2023, DPO) aims to forego both the reward modeling and online RL procedure. Rewriting (1) with a decomposition of the KL term into:
@@
Meta-Optimization for Algorithm Discovery. The goal of meta-optimization (optimizing the optimization process) is to uncover novel learning algorithms using a data-driven process. Suppose that an algorithm uses an objective function $f^{\gamma}$ to train a model for $K$ iterations, where $\gamma$ denotes a set of meta-parameters . Meta-optimization searches for an objective that maximizes the expected downstream performance $\max_{{}_{\gamma}}\mathbb{E}[\eta(\pi_{K})|\texttt{train}(f^{\gamma})]$ where $\eta$ is a downstream performance metric. Unlike previous methods that rely on a predefined parameterization of $\gamma$ (e.g., a neural network (Hospedales et al., 2021) or domain-specific language (Alet et al., 2020)), we leverage LLMs to directly propose code-level objective functions in Python. This approach eliminates the need for a carefully-designed search space and utilizes the extensive knowledge embedded in the LLM for flexible selection and mutation.
@@ 3 LLM-Driven Objective Discovery
4:     Run unit tests to check validity of the candidate and resample if needed.
@@ 4.1 Discovery Task - Multi-turn Dialogue on MT-Bench
Figure 3: Examples of LLM Objective Discovery improvement across generations. The first and second run shown left and right respectively.
@@ 4.2 Discovery Results
After evaluating approximately $100$ objective functions, we cataloged the best-performing ones in Table 1. We tabulate the high-level objective forms here and provide the full objective loss functions and their associated code in Appendix E. Moreover, we also plot the best performing sub-task evaluations in Figure 4.
@@
LRML (DiscoPOP) Log Ratio Modulated Loss $(1-\sigma(\rho/\tau))\cdot f_{dpo}(\beta\rho)+\sigma(\rho/\tau)\cdot f_{exp}(\beta\rho)$ 7.916
@@ 5 Held-Out Evaluations
We provide the Alpaca Eval 2.0 results in Table 2. As reference policies, we used GPT-4 for absolute comparison and the SFT-trained model for relative comparison. We observe that the discovered LRML (DiscoPOP), PADLL, and AQFL functions outperform the baselines and other discovered losses on the normal and length-controlled win rates. The differences in scores among these top performing losses are not significant, except for the LC win rate against the SFT reference model, where DiscoPOP performs best.
@@ 5.2 Summarization (TL;DR)
Table 3: TL;DR - Held Out Summarization Task Win rate of various preference optimization functions in the summarization task evaluated with the Alpaca Eval 2.0 calculations, against a subset of the test set (694 samples). The baseline outputs are the human generated preferences, and the model after SFT (see Appendix C for details). Note that the standard error in the LC win-rate has been rounded down because of values $<0.001$. We have highlighted the scores with means overlapping the standard error of the best score.
@@ 5.3 Positive sentiment generation (IMDb)
We provide results of models with converging $\beta$ values in Figure 5 for LRML compared against DPO and SLiC, displaying the model rewards against the KL-Divergence to the reference model. In Figure 5(a), the LRML-trained text generator outperforms the DPO model in terms of rewards and KL-divergence with low $\beta$ values (0.025, 0.05, 0.1). At higher $\beta$ values (0.5 and 1.0) both methods show trends of increased KL-Divergence and lower rewards, but generally LRML maintains a higher reward than DPO. In Figure 5(b), we note that LRML slightly outperforms DPO, SLiC, AQFL, and PADLL at $\beta\in\{0.05,0.1\}$ in terms of reward. For larger $\beta$ values (0.5 and 1.0), LRML shows similar trends of increased KL-Divergence and rewards like the other objective functions. A more detailed comparison between the individual discovered losses and the baselines can be found in Appendix Figure 8.
@@ 6.1 Log Ratio Modulated Loss (DiscoPOP)
$f_{lrml}(\beta\rho)$ $=(1-\sigma(\rho/\tau))\cdot f_{dpo}(\beta\rho)+\sigma(\rho/\tau)\cdot f_{exp}(\beta\rho)$ (4)
$=(1-\sigma(\rho/\tau))\cdot\text{log}(1+\text{exp}(-\beta\rho))+\sigma(\rho/\tau)\cdot\text{exp}(-\beta\rho)$ (5)
@@ 6.2 Limitations of DiscoPOP
In Figure 9 and Figure 10 of the Appendix, we plot the LRML objective function for $\beta\in\{0.01,0.025,0.05,0.1,0.25,0.5,1,2.5,5\}$ against DPO. Notably, when $\beta$ is high, the DiscoPOP objective function takes the form of the DPO log sigmoid loss. During training on $\beta=0.01$, we observed that DiscoPOP gets stuck in generating negative reviews. We hypothesize that it is because the loss is stuck in the local-minima to the left with negative difference of log-ratios. While training with $\beta\in{2.5,5.0}$ we observed that the model collapsed after a sharp spike in the loss and subsequently having loss value 0 and NaN outputs. This is potentially due to large gradient in the non-convex part, which could potentially be amended with gradient clipping.
@@ 7 Related Work
Evolution and Search with Large Language Models. LLMs provide a fast and automated way to create multiple candidate solutions for a problem stated in natural language (Song et al., 2024). This makes them powerful tools for driving population-based search procedures such as evolutionary meta-discovery. Various recent works have applied this approach to coding problems (Romera-Paredes et al., 2024), neural architecture search (Chen et al., 2024a), virtual robotic design settings (Lehman et al., 2023), and reward functions (Ma et al., 2023; Yu et al., 2023). Finally, recently LLMs have shown to be capable of acting as recombination operators for black-box optimization with Evolution Strategies (Lange et al., 2024) and for Quality-Diversity approaches (Lim et al., 2024).
Automated Discovery for Machine Learning. There are many other approaches to automating the discovery of generalizable machine learning algorithms. Some prior works explore the space of ML functions using genetic algorithms and a hand-crafted domain-specific language for reinforcement learning algorithms (Co-Reyes et al., 2021), curiosity algorithms (Alet et al., 2020), and optimizers (Chen et al., 2024b). Other works instead parameterize a transferrable objective function using neural networks and optimize them with evolution strategies. For example, Lu et al. (2022); Jackson et al. (2024); Houthooft et al. (2018); Alfano et al. (2024) evolve policy optimization objectives, Metz et al. (2022) evolves neural network optimizers, and Lange et al. (2023b, a) evolve blackbox optimizers.
@@ 8 Conclusion
Summary. In this paper, we proposed and used LLM-driven objective discovery to generate novel offline preference optimization algorithms. Specifically, we were able to discover high-performing preference optimization losses that achieve strog performance across held-out evaluation tasks, with the highest performing providing new insights into what an optimal objective may need to possess, such as being a blend of logistic and exponential losses, and possibly be non-convex.
Limitations & Future work. There are multiple limitations of our current approach. First, we have only scratched the surface of how to generate LLM objective proposals most effectively. Initial exploratory experiments using techniques such as temperature sampling or worst-to-best performance sorting in the context did not yield significant improvements. But one could imagine leveraging more information about the training runs and automatically tuning instruction prompt templates. E.g. by providing entire learning curve plots to a Visual Language Model (see Figure 12) or by meta-meta-optimizing (Lu et al., 2023) the LLM prompt. Second, the highest-performing loss re-purposed $\beta$ in the traditional sense, making it affect the functional behavior as well as the KL penalty of the model with respect to the base model. This motivates future work to study different forms, with perhaps multiple floating point parameters in the form, that each could be tuned separately. Although we provided an initial analysis sweep over this one single parameter and observed some instances of the functional behavior leading to instability of training the model, a further multi-parameter analysis, reformulating the objective, would be beneficial for future work. Finally, our work uses closed-source models (GPT-4) to generate code, which limits reproducibility and is costly to run. Future work could use the produced models themselves to generate code, resulting in code-level self-improvement.
@@
This work was supported by Azure sponsorship credits granted by Microsoft’s AI for Good Research Lab and by the Microsoft’s Accelerate Foundation Models Academic Research initiative. The hardware used for training was sponsored by GoodAI. SH is funded by AstraZeneca. CF is funded by Canon Medical. AJC is funded by a Microsoft Research and EPSRC ICASE scholarship award. The code can also be accessed at https://github.com/samholt/DiscoPOP.

### v2

@@ Abstract
Offline preference optimization is a key method for enhancing and controlling the quality of Large Language Model (LLM) outputs. Typically, preference optimization is approached as an offline supervised learning task using manually crafted convex loss functions. While these methods are based on theoretical insights, they are inherently constrained by human creativity, so the large search space of possible loss functions remains under-explored. We address this by performing LLM-driven objective discovery to automatically discover new state-of-the-art preference optimization algorithms without (expert) human intervention. Specifically, we iteratively prompt an LLM to propose and implement new preference optimization loss functions based on previously evaluated performance metrics. This process leads to the discovery of previously unknown and performant preference optimization algorithms. The best performing of these we call Discovered Preference Optimization (DiscoPOP) [^1] Code: https://github.com/luchris429/DiscoPOP., a novel algorithm that adaptively blends logistic and exponential losses. Experiments demonstrate the state-of-the-art performance of DiscoPOP and its successful transfer to held-out tasks.
@@ 1 Introduction
In this work, we aim to address these limitations by performing LLM-driven discovery to automatically generate new state-of-the-art preference optimization algorithms without continual expert human intervention in the development process. While previous works (Ma et al., 2023; Yu et al., 2023) have used LLMs to design environment-specific RL reward functions, we discover general-purpose objective functions which can be used across various preference optimization tasks. More specifically, we iteratively prompt an LLM to propose new preference optimization loss functions and evaluate them, with the previously proposed loss functions and their task performance metric (in our case, MT-Bench scores (Zheng et al., 2024)) as in-context examples. After performing this automatic discovery process, we catalogue high-performing loss functions and introduce a particularly strong one we call Discovered Preference Optimization (DiscoPOP), a new algorithm. To ensure robustness beyond MT-Bench, we validate DiscoPOP using AlapacaEval 2.0 (Dubois et al., 2024), showing an improvement in win rates against GPT-4 from DPO $(11.23\%\rightarrow 13.21\%)$. Additionally, in separate, held-out, tasks such as summarization and controlled generation, models trained with the DiscoPOP loss outperform or perform competitively with existing preference optimization algorithms.
@@
Figure 1: Left. Conceptual illustration of LLM-driven discovery of objective functions. We prompt an LLM to output new code-level implementations of offline preference optimization losses $\mathbb{E}_{(y_{w},y_{l},x)\sim\mathcal{D}}\left[f\left(\beta\rho\right)\right]$ as a function of the policy ($\pi_{\theta}$) and reference model’s ($\pi_{\text{ref}}$) likelihoods of the chosen ($y_{w}$) and rejected ($y_{l}$) completions. Afterwards, we run an inner loop training procedure and evaluate the resulting model on MT-Bench. The corresponding performance is fed back to the language model, and we query it for the next candidate. Right. Performance of discovered objective functions on Alpaca Eval.
@@ 2 Background
Preference Optimization. Consider a pre-trained language model policy $\pi_{\theta}$ and a dataset $\mathcal{D}=\{(x^{i},y_{w}^{i},y_{l}^{i})\}_{i=1}^{N}$ consisting of prompts $x$ and preference-ranked completions $y_{w}$ and $y_{l}$. In this dataset, a human rater prefers $y_{w}$ over $y_{l}$, denoted as $y_{w}\succ y_{l}$. The task is to align $\pi_{\theta}$ with the human values implicit in these preferences. Canonically, this has been achieved through reinforcement learning from human feedback (Christiano et al., 2017, RLHF), an approach that proceeds in two phases: First, a reward modelling stage that learns a parameterized reward model $r_{\phi}$. By assuming a Bradley-Terry model (Bradley and Terry, 1952) of preferences, the probability of the data can be expressed as $P(y_{w}\succ y_{l})=\exp r_{\phi}({y_{w},x})/(\exp r_{\phi}({y_{w},x})+\exp r_{\phi}({y_{l},x}))$, and subsequently simply optimized over $\phi$ through the maximum likelihood principle. The second stage of policy optimization employs a reinforcement learning algorithm to train the language model against the learned reward. Usually, a KL penalty is introduced between the model and the pre-RL reference policy $\pi_{ref}$ (Jaques et al., 2019; Stiennon et al., 2020) to prevent over-optimization and straying too far from the original policy, resulting in the final objective:
@@
Despite success in frontier models (Anthropic, 2023; Gemini-Team, 2023), deep RL has many implementations (Engstrom et al., 2019) and training challenges (Sutton, 1984; Razin et al., 2023) that hinder its adoption. To simplify the whole process, direct preference optimization (Rafailov et al., 2023, DPO) aims to forego both the reward modelling and online RL procedure. Rewriting (1) with a decomposition of the KL term into:
@@
Meta-Optimization for Algorithm Discovery. The goal of meta-optimization (optimizing the optimization process) is to uncover novel learning algorithms using a data-driven process. Suppose that an algorithm uses an objective function $f^{\gamma}$ to train a model for $K$ iterations, where $\gamma$ denotes a set of meta-parameters. Meta-optimization searches for an objective that maximizes the expected downstream performance $\max_{{}_{\gamma}}\mathbb{E}[\eta(\pi_{K})|\texttt{train}(f^{\gamma})]$ where $\eta$ is a downstream performance metric. Unlike previous methods that rely on a predefined parameterization of $\gamma$ (e.g., a neural network (Hospedales et al., 2021) or domain-specific language (Alet et al., 2020)), we leverage LLMs to directly propose code-level objective functions in Python. This approach eliminates the need for a carefully designed search space and utilizes the extensive knowledge embedded in the LLM for flexible selection and mutation.
@@ 3 LLM-Driven Objective Discovery
4:     Run unit tests to check the validity of the candidate and resample if needed.
@@ 4.1 Discovery Task - Multi-turn Dialogue on MT-Bench
Figure 3: Examples of LLM Objective Discovery improvement across generations. The first and second runs are shown left and right respectively.
@@ 4.2 Discovery Results
After evaluating approximately $100$ objective functions, we catalogued the best-performing ones in Table 1. We tabulate the high-level objective forms here and provide the full objective loss functions and their associated code in Appendix E. Moreover, we also plot the best performing sub-task evaluations in Figure 4.
@@
LRML (DiscoPOP) Log Ratio Modulated Loss $(1-\sigma(\beta\rho/\tau))\cdot f_{dpo}(\beta\rho)+\sigma(\beta\rho/\tau)\cdot f_{exp}(\beta\rho)$ 7.916
@@ 5 Held-Out Evaluations
We provide the Alpaca Eval 2.0 results in Table 2. As reference policies, we used GPT-4 for absolute comparison and the SFT-trained model for relative comparison. We observe that the discovered LRML (DiscoPOP), PADLL, and AQFL functions outperform the baselines and other discovered losses on the normal and length-controlled win rates. The differences in scores among these top-performing losses are not significant, except for the LC win rate against the SFT reference model, where DiscoPOP performs best.
@@ 5.2 Summarization (TL;DR)
Table 3: TL;DR - Held Out Summarization Task Win-rate of various preference optimization functions in the summarization task was evaluated with the Alpaca Eval 2.0 calculations, against a subset of the test set (694 samples). The baseline outputs are the human-generated preferences, and the model after SFT (see Appendix C for details). Note that the standard error in the LC win-rate has been rounded down because of values $<0.001$. We have highlighted the scores with means overlapping the standard error of the best score.
@@ 5.3 Positive sentiment generation (IMDb)
We provide results of models with converging $\beta$ values in Figure 5 for LRML compared against DPO and SLiC, displaying the model rewards against the KL-Divergence to the reference model. In Figure 5(a), the LRML-trained text generator outperforms the DPO model in terms of rewards and KL-divergence with low $\beta$ values (0.025, 0.05, 0.1). At higher $\beta$ values (0.5 and 1.0) both methods show trends of increased KL-Divergence and lower rewards, but generally, LRML maintains a higher reward than DPO. In Figure 5(b), we note that LRML slightly outperforms DPO, SLiC, AQFL, and PADLL at $\beta\in\{0.05,0.1\}$ in terms of reward. For larger $\beta$ values (0.5 and 1.0), LRML shows similar trends of increased KL-Divergence and rewards like the other objective functions. A more detailed comparison between the individual discovered losses and the baselines can be found in Appendix Figure 8.
@@ 6.1 Log Ratio Modulated Loss (DiscoPOP)
$f_{lrml}(\beta\rho)$ $=(\sigma(\beta\rho/\tau)-1)\cdot f_{dpo}(\beta\rho)+\sigma(\beta\rho/\tau)\cdot f_{exp}(\beta\rho)$ (4)
$=(1-\sigma(\beta\rho/\tau))\cdot\text{log}(1+\text{exp}(-\beta\rho))+\sigma(\beta\rho/\tau)\cdot\text{exp}(-\beta\rho)$ (5)
@@ 6.2 Limitations of DiscoPOP
In Figure 9 and Figure 10 of the Appendix, we plot the LRML objective function for $\beta\in\{0.01,0.025,0.05,0.1,0.25,0.5,1,2.5,5\}$ against DPO. When $\beta$ is high, the DiscoPOP objective function takes the form of the DPO log sigmoid loss. During training on $\beta=0.01$, we observed that DiscoPOP gets stuck in generating predominantly negative reviews (resulting in a reward score of $\sim 0.15$). We hypothesize that the loss is stuck in the local minima to the left with a negative difference in log ratios. While training with $\beta\in\{2.5,5.0\}$ we observed that the model collapsed after a sharp spike in the loss and subsequently having loss value 0 and NaN outputs. This is potentially due to a large gradient in the non-convex part, which could be amended with gradient clipping.
@@ 7 Related Work
Evolution and Search with Large Language Models. LLMs provide a fast and automated way to create multiple candidate solutions for a problem stated in natural language (Song et al., 2024), which makes them powerful tools for driving population-based search procedures. Various recent works have applied this approach to coding problems (Romera-Paredes et al., 2024), neural architecture search (Chen et al., 2024a), virtual robotic design settings (Lehman et al., 2023), reward functions (Ma et al., 2023; Yu et al., 2023), and algorithm heuristics (Liu et al., 2024). Finally, recently LLMs have shown to be capable of acting as recombination operators for black-box optimization with Evolution Strategies (Lange et al., 2024) and for Quality-Diversity approaches (Lim et al., 2024).
Automated Discovery for Machine Learning. There are many other approaches to automating the discovery of generalizable machine learning algorithms. Some prior works explore the space of ML functions using genetic algorithms and a hand-crafted domain-specific language for reinforcement learning algorithms (Co-Reyes et al., 2021), curiosity algorithms (Alet et al., 2020), and optimizers (Chen et al., 2024b). Other works instead parameterize a transferrable objective function using neural networks and optimize them with evolution strategies or meta-gradients. For example, Lu et al. (2022); Jackson et al. (2024b); Houthooft et al. (2018); Alfano et al. (2024); Kirsch et al. (2019); Oh et al. (2020); Jackson et al. (2024a) discover policy optimization objectives, Metz et al. (2022) evolves neural network optimizers, and Lange et al. (2023b, a) evolve blackbox optimizers.
@@ 8 Conclusion
Summary. In this paper, we proposed and used LLM-driven objective discovery to generate novel offline preference optimization algorithms. Specifically, we were able to discover high-performing preference optimization losses that achieve strong performance across held-out evaluation tasks, with the highest performing providing new insights into what an optimal objective may need to possess, such as being a blend of logistic and exponential losses, and possibly being non-convex.
Limitations & Future work. There are multiple limitations to our current approach. First, we have only scratched the surface of how to generate LLM objective proposals most effectively. Initial exploratory experiments using techniques such as temperature sampling or worst-to-best performance sorting in the context did not yield significant improvements. But one could imagine leveraging more information about the training runs and automatically tuning instruction prompt templates. E.g. by providing entire learning curve plots to a Visual Language Model (see Figure 12) or by meta-meta-optimizing (Lu et al., 2023) the LLM prompt. Second, the highest-performing loss re-purposed $\beta$ in the traditional sense, making it affect the functional behavior as well as the KL penalty of the model with respect to the base model. This motivates future work to study different forms, with perhaps multiple floating point parameters in the form, that each could be tuned separately. Although we provided an initial analysis sweep over this one single parameter and observed some instances of the functional behavior leading to instability of training the model, a further multi-parameter analysis, reformulating the objective, would be beneficial for future work. Finally, our work uses closed-source models (GPT-4) to generate code, which limits reproducibility and is costly to run. Future work could use the produced models themselves to generate code, resulting in code-level self-improvement.
@@
This work was supported by Azure sponsorship credits granted by Microsoft’s AI for Good Research Lab and by Microsoft’s Accelerate Foundation Models Academic Research initiative. The hardware used for training was sponsored by GoodAI. SH is funded by AstraZeneca. CF is funded by Canon Medical. AJC is funded by a Microsoft Research and EPSRC ICASE scholarship award. The code can also be accessed at https://github.com/samholt/DiscoPOP.